Find the frequency of 1 Mev photon. Given wavelength of a 1 Kev photon is 1.24*10-9 ?
 
Pls ans asap

Grade:12th pass

1 Answers

Arun
5 years ago
Energy E = hν = hc/λ = 1 keV  -----------(1)
Wavelength λ = 1.24 x 10-9 m  -----------(2)
From (1) and (2), we get
hc = λ x 1 keV = 1.24 x 10-9 m x 1 keV = 1.24 x 10-9 m. 103 eV = 1.24 x 10-6 eV.m

If E = hc/λ = 1 MeV
  1/λ = E/hc = 1 x 106 eV/(1.24 x 10-6 eV.m)
                  = 0.81 x 1012/m
Frequency ν = c/λ = (3 x 108 m/s) x 0.81 x 1012/m
                                = 2.43x 1020/s
                                 = 2.43 x 1020 Hz.